OpenStudy (notamathgenius):

In a​ triangle, the measure of the first angle is twice the measure of the second angle. The measure of the third angle is 76° more than the measure of the second angle. Use the fact that the sum of the measures of the three angles of a triangle is 180° to find the measure of each angle.

OpenStudy (notamathgenius):

@Benji

OpenStudy (benji):

Alright  everything make sense?

OpenStudy (notamathgenius):

how do i write that into formula? i did it this way idk if thats right tho x+76+2x+x=180

OpenStudy (benji):

that is right, now you can combine like terms and solve for x

OpenStudy (notamathgenius):

x+x= 2x correct?

OpenStudy (benji):

How did you get that? Where did the 180 and 76 go?

OpenStudy (notamathgenius):

nono, im saying to combine like terms

OpenStudy (benji):

oh yes that is right

OpenStudy (notamathgenius):

ok so after all the combining i got, 4x+76=180

OpenStudy (benji):

Yep

OpenStudy (notamathgenius):